OpenStudy (bbb911):

[7 (x-3) + 4] - [6(3x-2) + x] =

OpenStudy (mayankdevnani):

@bbb911 first solve \[\huge 7(x-3)+4\]

OpenStudy (jhannybean):

-_-

OpenStudy (bbb911):

7x-17

OpenStudy (mayankdevnani):

correct

OpenStudy (mayankdevnani):

then solve \[\huge 6(3x-2)+x\]

OpenStudy (jhannybean):

there is a negative.

OpenStudy (mayankdevnani):

no we solve it later

OpenStudy (mayankdevnani):

@bbb911

OpenStudy (anonymous):

=[7x-21+4] - [18x-12 +x] =[7x-17] - [19x -12] =-12x -5 =-(12x+5)

OpenStudy (bbb911):

so i don't include the negtive

OpenStudy (anonymous):

piece of cake.

OpenStudy (mayankdevnani):

@bbb911 negative...we solve it later

OpenStudy (bbb911):

kk :

OpenStudy (mayankdevnani):

so solve

OpenStudy (bbb911):

im still gonna solve it

OpenStudy (mayankdevnani):

k

OpenStudy (bbb911):

since its 19x-12

OpenStudy (bbb911):

then you minus that to 1x-17 right?

OpenStudy (mayankdevnani):

correct

OpenStudy (mayankdevnani):

then, \[\huge 7x-17-(19x-12)\]

OpenStudy (mayankdevnani):

can you solve it??? @bbb911

OpenStudy (bbb911):

ill try

OpenStudy (mayankdevnani):

please

OpenStudy (jhannybean):

( 7x - 19x ) -17 + 12 = 0

OpenStudy (bbb911):

al i will do is subtract right?

OpenStudy (mayankdevnani):

@tanjeetsarkar96 and for your kind information, please don't give direct answers!!!

OpenStudy (mayankdevnani):

@bbb911 first change the sign

OpenStudy (jhannybean):

@mayankdevnani ask her togroup them, it makes the equation easier tosolve. Then ask her what 7-19 is. Then ask her what -17 + 12 is. Then have her solve for x. lol.

OpenStudy (mayankdevnani):